There is a competition at a local movie theater for free movie tickets. You must guess all four employees ages given at few clues. The first clue is that when added together, their ages total is 106 years. Kirk is twice 10 years less than the managers age, brian is 12 years younger than twice the manager’s age, and matt is 6 years older than half the manager’s age. What are all four of their ages?
Found 3 solutions by Boreal, addingup, MathTherapy:
Answer by Boreal(15235)

manager's age is x
kirk is 2(x-10)
brian is 2x-12
matt is (1/2)x+6
these add to 106
so 2x-20+2x-12+(1/2)x+6+x=106
5.5x=132
x=24 Manager
Kirk is 28
Brian 36
Matt 18

Answer by addingup(3677)

k+M+b+m = 106 (M = manager, m = Matt)
k = M - 2(10) = M - 20
b = 2M - 12
m = 1/2M+6
.
M - 20 + 2M - 12 + 1/2M + 6 = 106
3.5M - 26 = 106
3.5M = 132
M = 37.7
~~~~~~~~~~~~~~~~~
Manager = 37.7
Kirk = 17.7
Brian = 63.4
Matt = 24.9

Answer by MathTherapy(10555)


There is a competition at a local movie theater for free movie tickets. You must guess all four employees ages given at few clues. The first clue is that when added together, their ages total is 106 years. Kirk is twice 10 years less than the managers age, brian is 12 years younger than twice the manager’s age, and matt is 6 years older than half the manager’s age. What are all four of their ages?
@ADDINGUP's answers DON'T MAKE SENSE!
DON'T use them!